# Direct Imaging of Hydrogen-Driven Dislocation and Strain Field Evolution in a Stainless Steel Grain

### v2

@@ 3 Conclusion
Our findings showcase the unique capability of in situ BCDI to monitor the evolution of nanoscale deformation and defects in a grain exposed to hydrogen under bulk conditions. This advancement is particularly timely as third-generation synchrotrons are being upgraded to fourth-generation facilities, offering orders of magnitude greater coherent flux for BCDI experiments that will enhance spatial and temporal resolution. This could enable exciting possibilities for exploring hydrogen induced evolution in adjacent grains or extended grains using Bragg ptychography [62] In situ BCDI is applicable to most alloy systems and perfectly complements higher-resolution electron microscopy techniques, computational models, and atomistic simulations to shed light on HE degradation pathways. Understanding these internal mechanisms is key to designing alloys with enhanced resistance to hydrogen-induced degradation, such as engineering micro- and nano-structures to mitigate dislocation movement or vacancy formation. These hydrogen-resistant materials are urgently needed to enable the green hydrogen economy.
4 Methods
@@ 4.1 Sample preparation
The SS sample was produced using high pressure torsion (HPT) to produce sub-micron grain sizes appropriate for BCDI. A sheet of 316 A4 austenitic stainless steel (composition shown in Table LABEL:supp-table:composition), 1 mm thick, was obtained from RS Components (stock number 264-7241) and cut into a 5 mm diameter disk. The disk then went through HPT to refine the grain microstructure. HPT was performed in a quasi-constrained set-up [50] on a Zwick Roell Z100 Materials Testing machine. Using 5 mm diameter anvils, a compressive force of 80 kN was applied to the face of the disk and maintained for 30 turns using a rotational speed of 5 ° $\mathrm{s^{-1}}$. The final disk thickness after HPT was 300 $\mu$m. The disk was then annealed in vacuum ($5\times 10^{-6}$ mbar) at 700 °C for 1 hour, ramping up and down at 4 °C $\mathrm{min}^{-1}$. Following annealing, the disk was sequentially ground with 800, 1200, and 2500 grit SiC paper, followed by polishing in 3 $\mu$m and 1 $\mu$m diamond suspension.
Powder X-ray diffraction (Figure LABEL:supp-fig:powder_XRD) shows that after HPT, the sample will be predominately composed of deformation-induced martensite. The heat treatment removes macroscale residual stresses and the sample undergoes recovery and recrystallization to form austenite [63, 64]. The measured austenite grain was likely formed during heat treatment due to the relatively low misorientation and very low observed strain (relative to the average lattice parameter of the grain) near the surface of the grain. We note that carbides could have formed during the annealing, which could attract solutes that have become mobile due to hydrogen.
@@ 4.2 BCDI measurements
BCDI relies on inverting far-field, oversampled [65], 3D coherent X-ray diffraction patterns (CXDPs) for a particular $hkl$ reflection on a finite crystalline material. This is done using phase retrieval algorithms [49] to obtain the grain morphology and associated phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$. Dislocation line positions can be readily identified from the positions of phase vortex loci [43], thus making BCDI a valuable tool to study in situ or in operando defect evolution [43, 44, 45, 46, 66]. The strain fields associated with the dislocations can be obtained from the phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$, since $\mathbf{\psi}_{hkl}(\mathbf{r})$ is a projection of the lattice displacement field, $\mathbf{u}_{hkl}(\mathbf{r})$, onto the scattering vector, $\mathbf{Q}_{hkl}$ [67].
BCDI was performed at beamline 34-ID-C at the Advanced Photon Source (APS) at Argonne National Laboratory, USA. An in situ confocal microscope was used to position the X-ray beam within the disk. The $111$ Bragg peak ($34.72\theta=$ 34.7°) was measured for multiple grains to screen for the best candidate grain. Although each candidate’s position perpendicular to the X-ray beam can be easily aligned, the position of the candidate along the beam needs to be determined to ensure the grain is on the axis of rotation. We used an approach presented by Shabalin et al. [68] to accomplish this.
For all BCDI measurements, the grain was illuminated using a 10 keV ($\lambda$ = 0.124 nm) coherent X-ray beam, with a bandwidth of $\Delta\lambda/\lambda\approx 10^{-4}$ from a Si$(111)$ monochromator. The X-ray beam was focused to a size of 810 nm × 860 nm (h × v, full width at half-maximum) using Kirkpatrick–Baez (KB) mirrors. Beam defining slits were used to select the coherent portion of the beam at the entrance to the KB mirrors. CXDPs were collected on a 256 × 256 pixel module of a 512 × 512 pixel Timepix area detector (Amsterdam Scientific Instruments) with a GaAs sensor and pixel size of $55\mathrm{\ \mu m}\times 55\mathrm{\ \mu m}$ positioned at 1.4 m from the sample to ensure oversampling. The peak of the CXDP was positioned at the center of the detector module before data collection. CXDPs were recorded by rotating the crystal through an angular range of 0.5° about the peak and recording an image every 0.005° with 0.5 s exposure time and 10 accumulations at each angle.
@@ 4.3 Hydrogen charging
The disk was attached to the electrochemical flow cell using chemically resistant insulating tape, with the bottom of the disk in contact with a Pt wire, thereby forming the working electrode (WE). The counter electrode (CE) was another Pt wire coiled into the well, or moat, of the electrochemical cell. The cell was then sealed using an O-ring and a thin kapton film. The flow cell was fixed to the sample stage using a Thorlabs 1X1 kinematic mount and connected to a Cole-Parmer Masterflex peristaltic pump and a SP-300 Biologic potentiostat. See Supporting Information and Figure LABEL:supp-fig:flowcell for further details about the electrochemical flow cell.
A 1 L hydrogen charging solution was prepared using 4 g of NaOH (0.1 mol $\mathrm{L}^{-1}$) and 5 g thiourea (0.07 mol $\mathrm{L}^{-1}$) dissolved in deionized water, yielding a pH of 13. After an appropriate grain was found, the hydrogen charging solution was continuously pumped into the cell at 5 mL $\mathrm{min}^{-1}$. Repeated $111$ CXDPs from the grain were measured in solution, without hydrogen charging, over three hours to verify that there were no X-ray effects on the sample. Furthermore, any X-ray induced changes to the strain field are negligible (Supporting Text). Next, hydrogen charging was performed using chronopotentiometry with a two electrode setup, keeping the current fixed at 0.2 mA (0.5 mA $\mathrm{cm^{-2}}$). During the charging, two repeated BCDI measurements for the $111$ reflection were measured approximately every 40 minutes, with each scan requiring approximately 10 minutes to complete. We noticed bubbles slowly form during charging.
@@ 4.5 BCDI strain calculations
$\varepsilon_{hkl}\mathbf{(r)}=\frac{\partial\mathbf{u}_{hkl}(\mathbf{r})}{\partial x_{hkl}}=\nabla\mathbf{\psi}_{hkl}(\mathbf{r})\cdot\frac{\mathbf{Q}_{hkl}}{|\mathbf{Q}_{hkl}|^{2}}.$ (3)
@@
$e_{hkl}=\frac{a_{0,t}-a_{0,\mathrm{ref.}}}{a_{0,\mathrm{ref.}}},$ (4)
where $a_{0,\mathrm{ref.}}$ is the reference lattice parameter, determined at -3.4 h. $a_{0}$ was calculated using Bragg’s law and the $2\theta$ angle determined from the angular position of the diffractometer.
@@ 4.6 Hydrogen concentration
$\varepsilon_{\mathrm{vol}}=\sum_{A}n^{(A)}\Omega^{(A)},$ (5)
where $n^{(A)}$ is the defect concentration for defect type $A$, and $\Omega^{(A)}$ is the relaxation volume for the specific defect type. The FCC metal lattice has two interstitial sites available for accommodating hydrogen atoms: octahedral (O) and tetrahedral (T) sites. In transition metals with FCC lattice, dissolved hydrogen atoms preferentially occupy the O site with larger free space than the T site [26]. We used a relaxation volume of $0.200\pm 0.005$ for interstitial hydrogen in austenitic stainless steels [51]. For our calculation, we assumed this value is similar for FCC 316 SS, and that all the interstitial hydrogen atoms reside in the O sites.
The volumetric strain relative to the initial volume at -3.4 h was calculated using the homogeneous strain, $e_{hkl}$ (Equation 4),
$\varepsilon_{\mathrm{vol}}=(1+e_{hkl})^{3}-1.$ (6)
@@
$n^{(A)}=\frac{(1+e_{hkl})^{3}-1}{\Omega^{(A)}}.$ (7)
The hydrogen concentrations presented in Figure 1b are reasonable when compared to values obtained from charging hydrogen into 316L SS [74].
@@ 4.7 Pearson correlation coefficient
The Pearson correlation coefficient, $r$, between two images was computed using Equation 8:
@@
where $x$ and $y$ were the Bragg peaks being compared, $x_{n}$ and $y_{n}$ were the values for a single voxel, and $\bar{x}$ and $\bar{y}$ were the means of each array.
@@ 4.8 Dislocation position identification
We define the dislocation line as the spatial positions of the locus of the phase vortices, referred to as nodes. Each dislocation is composed of many nodes, joined by edges based on the MATLAB graph object. The nodes were determined automatically by integrating the derivatives of the complex exponential of the phase $(e^{i\psi_{hkl}(\mathbf{r})})$ [75] and then selecting the maximum value as the dislocation node position if the value exceeded $0.75\pi$. The dislocation lines were found not to overlap perfectly in the reconstructions, which we attribute to noise and the limited spatial resolution. To increase overlap, the reconstructions were subpixel-shifted [76], on the order of a few pixels, or up to $\sim 30$ nm. The dislocation lines terminated on the surface of the grain.
@@ 4.9.1 Overview
Discrete dislocation dynamics (DDD) was used to simulate the motion of the dislocation structure and compare to the experimental result. The DDD simulations here are nodal, based on discrete straight segments [77, 78] presented by Yu et al. [79]. Dislocation motion is calculated by evaluating the velocity $\textbf{V}_{k}$ of node $k$ at position $\textbf{X}_{k}$ through a mobility law (shown in the following section), which describes mobility in various modes like glide, climb, and cross-slip.
@@ 4.9.2 FCC Mobility Law
$\Big{[}\frac{1}{2}\sum_{l}L^{kl}\boldsymbol{B}^{kl}\Big{]}^{-1}\boldsymbol{F}^{k}=\boldsymbol{V}^{k},$ (11)
where the sum is over all nodes $l$ connected to node $k$, and $\boldsymbol{F}^{k}$ is the nodal force determined by Equation 10.
@@ 4.9.3 Incorporating finite boundary conditions
To evaluate the corrective force term $\boldsymbol{\hat{F}}^{k}$ in Equation 10, the superposition principle is adopted. As described elsewhere [81, 82], the total stress strain and displacement fields are expressed as
@@
$\boldsymbol{u}$ $=\boldsymbol{\hat{u}}+\boldsymbol{\tilde{u}},$ (15)
@@ 4.9.4 Calculation Details
[…]
@@ 4.10 Dislocation strain field modeling
We assume that there are no externally applied stresses. The $\varepsilon_{111,\mathrm{model}}$ for the large dislocation was created using the dislocation node positions determined from BCDI, the Burgers vector (using the local lattice parameter, $a_{0,t}$), and $\nu=0.28$. Each dislocation node was joined by an edge, which was then connected to another point to form a dislocation triangle. The displacement field for this triangular dislocation loop was determined using the solution developed by Barnett [84, 85], which was then numerically differentiated to determine the lattice strain field, $\varepsilon_{\mathrm{model}}$,
$\varepsilon_{\mathrm{model}}=\begin{bmatrix}\varepsilon_{xx,\mathrm{model}}&\varepsilon_{xy,\mathrm{model}}&\varepsilon_{xz,\mathrm{model}}\\ \varepsilon_{yx,\mathrm{model}}&\varepsilon_{yy,\mathrm{model}}&\varepsilon_{yz,\mathrm{model}}\\ \varepsilon_{zx,\mathrm{model}}&\varepsilon_{zy,\mathrm{model}}&\varepsilon_{zz,\mathrm{model}}\end{bmatrix}.$ (16)
@@
$\varepsilon_{111,\mathrm{model}}=\begin{bmatrix}a&b&c\end{bmatrix}\begin{bmatrix}\varepsilon_{xx,\mathrm{model}}&\varepsilon_{xy,\mathrm{model}}&\varepsilon_{xz,\mathrm{model}}\\ \varepsilon_{yx,\mathrm{model}}&\varepsilon_{yy,\mathrm{model}}&\varepsilon_{yz,\mathrm{model}}\\ \varepsilon_{zx,\mathrm{model}}&\varepsilon_{zy,\mathrm{model}}&\varepsilon_{zz,\mathrm{model}}\end{bmatrix}\begin{bmatrix}a\\ b\\ c\end{bmatrix},$ (17)
@@
$\begin{bmatrix}a\\ b\\ c\end{bmatrix}=\begin{bmatrix}\mathbf{x^{\prime}}&\mathbf{y^{\prime}}&\mathbf{z^{\prime}}\end{bmatrix}^{-1}\begin{bmatrix}\mathbf{\hat{Q}_{hkl,\ sam}}\end{bmatrix},$ (18)
where $\mathbf{\hat{x}^{\prime}}$ is $\mathbf{\hat{b}_{sam}}$, $\mathbf{\hat{y}^{\prime}}$, is the dislocation loop plane normal vector (normalised) at -3.4 h, and $\mathbf{\hat{z}^{\prime}}=\mathbf{\hat{x}^{\prime}}\times\mathbf{\hat{y}^{\prime}}$. All vectors in Equation 18 are column vectors. The subscript, $\mathbf{sam}$, corresponds to the the sample coordinates relative to the standard unit vectors $\mathbf{\hat{x}}$, $\mathbf{\hat{y}}$, and $\mathbf{\hat{z}}$ axes of a 3D Cartesian coordinate system.

### v4

@@ 3 Conclusion
Our findings showcase the unique capability of in situ BCDI to monitor the evolution of nanoscale deformation and defects in a grain exposed to hydrogen under bulk conditions. This advancement is particularly timely as third-generation synchrotrons are being upgraded to fourth-generation facilities, offering orders of magnitude greater coherent flux for BCDI experiments that will enhance spatial and temporal resolution. This could enable exciting possibilities for exploring hydrogen-induced evolution in adjacent grains or extended grains using Bragg ptychography [72]. In situ BCDI is applicable to most alloy systems and perfectly complements higher-resolution electron microscopy techniques, computational models, and atomistic simulations to shed light on HE degradation pathways. Using the example of 316 SS, our results show the hydrogen-driven evolution of a single dislocation undergoing glide, followed by climb. We propose that glide is initiated by the diffusion of hydrogen into the lattice, where it reduces pinning strength and thus enables the dislocation to glide in-plane, reducing its line length and thus elastic energy. Our strain measurements also show that hydrogen shields the elastic strain field associated with the dislocation. With increasing hydrogen concentration, we observe rapid dislocation climb that is not expected at room temperature. This suggests that hydrogen alters the energy landscape, such that vacancies and interstitials become more readily available at the dislocation line, thus enabling climb motion. Understanding these internal mechanisms is key to designing alloys with enhanced resistance to hydrogen-induced degradation, such as engineering micro- and nano-structures to mitigate dislocation movement or vacancy formation [73]. These hydrogen-resistant materials are urgently needed to enable the green hydrogen economy to help achieve decarbonization targets.
4 Experimental Section
@@ 4.1 Sample preparation
[…]
Powder X-ray diffraction (Figure LABEL:supp-fig:powder_XRD, Supporting Information) showed that after HPT, the sample was predominately composed of deformation-induced martensite. The heat treatment removed macroscale residual stresses and the sample underwent recovery and recrystallization to form austenite [74, 75]. The measured austenite grain was likely formed during heat treatment due to the relatively low misorientation and very low observed strain (relative to the average lattice parameter of the grain) near the surface of the grain. It was noted that carbides could have formed during the annealing, which could attract solutes that have become mobile due to hydrogen.
@@ 4.2 BCDI measurements
BCDI relies on inverting far-field, oversampled [76], 3D coherent X-ray diffraction patterns (CXDPs) for a particular hklhkl reflection from a finite crystalline material. This was done using phase retrieval algorithms [52] to obtain the grain morphology and associated phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$. Dislocation line positions can be readily identified from the positions of phase vortex loci [46], thus making BCDI a valuable tool to study in situ or in operando defect evolution [46, 47, 48, 49, 77]. The strain fields associated with the dislocations can be obtained from the phase, $\mathbf{\psi}_{hkl}(\mathbf{r})$, since $\mathbf{\psi}_{hkl}(\mathbf{r})$ is a projection of the lattice displacement field, $\mathbf{u}_{hkl}(\mathbf{r})$, onto the scattering vector, $\mathbf{Q}_{hkl}$ [78].
BCDI was performed at beamline 34-ID-C at the Advanced Photon Source (APS) at Argonne National Laboratory, USA. An in situ confocal microscope was used to position the X-ray beam within the disk. The $111$ Bragg peak ($2\theta=34.7$°) was measured for multiple grains to screen for the best candidate grain. Although each candidate’s position perpendicular to the X-ray beam could be easily aligned, the position of the candidate along the beam needs to be determined to ensure the grain is on the axis of rotation. An approach presented by Shabalin et al. [79] was used to accomplish this.
For all BCDI measurements, the grain was illuminated using a 10 keV ($\lambda$ = 0.124 nm) coherent X-ray beam, with a bandwidth of $\Delta\lambda/\lambda\approx 10^{-4}$ from a Si(111)(111) monochromator. The X-ray beam was focused to a size of 810 nm × 860 nm (h × v, full width at half-maximum) using Kirkpatrick–Baez (KB) mirrors. Beam defining slits were used to select the coherent portion of the beam at the entrance to the KB mirrors. CXDPs were collected on a 256 × 256 pixel module of a 512 × 512 pixel Timepix area detector (Amsterdam Scientific Instruments) with a GaAs sensor and pixel size of $55\mathrm{\ \mu m}\times 55\mathrm{\ \mu m}$ positioned at 1.4 m from the sample to ensure oversampling. The peak of the CXDP was positioned at the center of the detector module before data collection. CXDPs were recorded by rotating the crystal through an angular range of 0.5° about the peak and recording an image every 0.005° with 0.5 s exposure time and 10 accumulations at each angle.
@@ 4.3 Hydrogen charging
The disk was attached to the electrochemical flow cell using chemically resistant insulating tape, with the bottom of the disk in contact with a Pt wire, thereby forming the working electrode (WE). The counter electrode (CE) was another Pt wire coiled into the well, or moat, of the electrochemical cell. The cell was then sealed using an O-ring and a thin kapton film. The flow cell was fixed to the sample stage using a Thorlabs 1X1 kinematic mount and connected to a Cole-Parmer Masterflex peristaltic pump and a SP-300 Biologic potentiostat. See Supporting Information and Figure LABEL:supp-fig:flowcell, Supporting Information for further details about the electrochemical flow cell.
A 1 L hydrogen charging solution was prepared using 4 g of NaOH (0.1 mol $\mathrm{L}^{-1}$) and 5 g thiourea (0.07 mol $\mathrm{L}^{-1}$) dissolved in deionized water, yielding a pH of 13. After an appropriate grain was found, the hydrogen charging solution was continuously pumped into the cell at 5 mL $\mathrm{min}^{-1}$. Repeated $111$ CXDPs from the grain were measured in solution, without hydrogen charging, over three hours to verify that there were no X-ray effects on the sample. Furthermore, any X-ray induced changes to the strain field are negligible (Supporting Information). Next, hydrogen charging was performed using chronopotentiometry with a two electrode setup, keeping the current fixed at 0.2 mA (0.5 mA $\mathrm{cm^{-2}}$). During the charging, two repeated BCDI measurements for the $111$ reflection were measured approximately every 40 min, with each scan requiring approximately 10 min to complete. Bubbles were noticed to slowly form during charging.
@@ 4.5 BCDI strain calculations
$\varepsilon_{hkl}\mathbf{(r)}=\frac{\partial\mathbf{u}_{hkl}(\mathbf{r})}{\partial x_{hkl}}=\nabla\mathbf{\psi}_{hkl}(\mathbf{r})\cdot\frac{\mathbf{Q}_{hkl}}{|\mathbf{Q}_{hkl}|^{2}}$ (3)
@@
$e_{hkl}=\frac{a_{0,\ t}-a_{0\mathrm{,\ ref.}}}{a_{0\mathrm{,\ ref.}}}$ (4)
where $a_{0\mathrm{,\ ref.}}$ is the reference lattice parameter, determined at –3.4 h. $a_{0}$ was calculated using Bragg’s law and the $2\theta$ angle determined from the angular position of the diffractometer.
@@ 4.6 Hydrogen concentration
$\varepsilon_{\mathrm{vol}}=\sum_{A}C^{(A)}\Omega_{R}^{(A)},$ (5)
where $C^{(A)}$ is the defect concentration for defect type $A$, and $\Omega_{R}^{(A)}$ is the relaxation volume for the specific defect type. The FCC metal lattice has two interstitial sites available for accommodating hydrogen atoms: octahedral (O) and tetrahedral (T) sites. In transition metals with FCC lattice, dissolved hydrogen atoms preferentially occupy the O site with larger free space than the T site [29]. $\Omega_{R}=0.200\pm 0.005$ was used for interstitial hydrogen in austenitic stainless steels [54]. For the calculation, this value was assumed to be similar for FCC 316 SS, and that all the interstitial hydrogen atoms reside in the O sites.
The volumetric strain relative to the initial volume at –3.4 h was calculated using the homogeneous strain, $e_{hkl}$ (Equation 4),
$\varepsilon_{\mathrm{vol}}=(1+e_{hkl})^{3}-1$ (6)
@@
$C=\frac{(1+e_{hkl})^{3}-1}{\Omega_{R}}$ (7)
The hydrogen concentrations presented in Figure 1b are reasonable when compared to values obtained from charging hydrogen into 316L $S$ [84].
@@ 4.7 Pearson correlation coefficient
The Pearson correlation coefficient, $r$, between two images was computed using Equation (8):
@@
where $x$ and $y$ are the Bragg peaks being compared, $x_{n}$ and $y_{n}$ are the values for a single voxel, and $\bar{x}$ and $\bar{y}$ are the means of each array.
@@ 4.8 Dislocation position identification
The dislocation line is identified as the spatial positions of the locus of the phase vortices, referred to as nodes. Each dislocation was composed of many nodes, joined by edges based on the MATLAB graph object. The nodes were determined automatically by integrating the derivatives of the complex exponential of the phase $(e^{i\psi_{hkl}(\mathbf{r})})$ [85] and then selecting the maximum value as the dislocation node position if the value exceeded $0.75\pi$. The dislocation lines were found not to overlap perfectly in the reconstructions, which were attributed to noise and the limited spatial resolution. To increase overlap, the reconstructions were subpixel-shifted [86], on the order of a few pixels, or up to $\approx 30$ nm. The dislocation lines terminated on the surface of the grain.
@@ 4.9.1 Overview
DDD was used to simulate the motion of the dislocation structure and compare to the experimental result. The DDD simulations here are nodal, based on discrete straight segments [87, 88] presented by Yu et al. [89]. Dislocation motion is calculated by evaluating the velocity $\textbf{V}_{k}$ of node $k$ at position $\textbf{X}_{k}$ through a mobility law (shown in the following section), which describes mobility in various modes like glide, climb, and cross-slip.
@@ 4.9.2 FCC Mobility Law
$\Big{[}\frac{1}{2}\sum_{l}L^{kl}\boldsymbol{B}^{kl}\Big{]}^{-1}\boldsymbol{F}^{k}=\boldsymbol{V}^{k}$ (11)
where the sum is over all nodes $l$ connected to node $k$, and $\boldsymbol{F}^{k}$ is the nodal force determined by Equation (10).
@@ 4.9.3 Incorporating finite boundary conditions
To evaluate the corrective force term $\boldsymbol{\hat{F}}^{k}$ in Equation (10), the superposition principle is adopted. As described elsewhere [91, 92], the total stress strain and displacement fields are expressed as
@@
$\boldsymbol{u}$ $=\boldsymbol{\hat{u}}+\boldsymbol{\tilde{u}}$ (15)
@@ 4.9.4 Calculation Details
The initial dislocation structure used in the simulation is identical to the initial dislocation structure measured by experiment before the start of hydrogen charging. The following parameters were used to represent $\gamma$-Fe: the lattice parameter, $a_{0}=3.602$ Å, the shear modulus, G=77G=77\ GPa, Poisson’s ratio, $\nu=0.28$, The dislocation structure was enclosed within a 0.6 $\mathrm{\mu}$m $\times$ 0.6 $\mathrm{\mu}$m $\times$ 0.6 $\mathrm{\mu}$m domain representing the experimentally observed grain. The finite element (FE) mesh was $20\times 20\times 20$. The drag coefficients for edge and screw dislocations were both 1.0, essentially making the mobility of each type of dislocation equivalent. In contrast, the drag coefficient for dislocation segments attempting to move in the climb direction was $10^{8}$, essentially confining the segments to the glide plane. The two end points were fixed for the entire simulation, while the internal nodes were free in all directions (though as discussed, they were restricted in the climb direction). No external mechanical load was applied to the domain.
@@ 4.10 Dislocation strain field modeling
It was assumed that there were no externally applied stresses. The $\varepsilon_{111,\ \mathrm{model}}$ for the large dislocation was created using the dislocation node positions determined from BCDI, the Burgers vector (using the local lattice parameter, $a_{0,\ t}$), and $\nu=0.28$. Each dislocation node was joined by an edge, which was then connected to another point to form a dislocation triangle. The displacement field for this triangular dislocation loop was determined using the solution developed by Barnett [94, 95], which was then numerically differentiated to determine the lattice strain field, $\varepsilon_{\mathrm{model}}$,
$\varepsilon_{\mathrm{model}}=\begin{bmatrix}\varepsilon_{xx,\ \mathrm{model}}&\varepsilon_{xy,\ \mathrm{model}}&\varepsilon_{xz,\ \mathrm{model}}\\ \varepsilon_{yx,\ \mathrm{model}}&\varepsilon_{yy,\ \mathrm{model}}&\varepsilon_{yz,\ \mathrm{model}}\\ \varepsilon_{zx,\ \mathrm{model}}&\varepsilon_{zy,\ \mathrm{model}}&\varepsilon_{zz,\ \mathrm{model}}\end{bmatrix}$ (16)
@@
$\varepsilon_{111,\ \mathrm{model}}=\begin{bmatrix}a&b&c\end{bmatrix}\begin{bmatrix}\varepsilon_{xx,\ \mathrm{model}}&\varepsilon_{xy,\ \mathrm{model}}&\varepsilon_{xz,\ \mathrm{model}}\\ \varepsilon_{yx,\ \mathrm{model}}&\varepsilon_{yy,\ \mathrm{model}}&\varepsilon_{yz,\ \mathrm{model}}\\ \varepsilon_{zx,\ \mathrm{model}}&\varepsilon_{zy,\ \mathrm{model}}&\varepsilon_{zz,\ \mathrm{model}}\end{bmatrix}\begin{bmatrix}a\\ b\\ c\end{bmatrix}$ (17)
@@
$\begin{bmatrix}a\\ b\\ c\end{bmatrix}=\begin{bmatrix}\mathbf{x^{\prime}}&\mathbf{y^{\prime}}&\mathbf{z^{\prime}}\end{bmatrix}^{-1}\begin{bmatrix}\mathbf{\hat{Q}_{hkl,\ sam}}\end{bmatrix}$ (18)
where $\mathbf{\hat{x}^{\prime}}$ is $\mathbf{\hat{b}_{sam}}$, $\mathbf{\hat{y}^{\prime}}$, is the dislocation loop plane normal vector (normalized) at –3.4 h, and $\mathbf{\hat{z}^{\prime}}=\mathbf{\hat{x}^{\prime}}\times\mathbf{\hat{y}^{\prime}}$. All vectors in Equation (18) are column vectors. The subscript, $\mathbf{sam}$, corresponds to the the sample coordinates relative to the standard unit vectors $\mathbf{\hat{x}}$, $\mathbf{\hat{y}}$, and $\mathbf{\hat{z}}$ axes of a 3D Cartesian coordinate system.